OpenStudy (anonymous):

1-1/4(2-(3-x)/(5)) ≤ (8-x)/(3)

OpenStudy (demonickittenlover):

o.o wow I have no clue im sorry wish I could help

hero (hero):

@L-Lawliet-L, Can you write it out using the draw feature?

hero (hero):

\[1 - \frac{1}{4\left(2-\frac{3-x}{5}\right)} \le \frac{8-x}{3}\]

hero (hero):

That's what I think you posted.

OpenStudy (anonymous):

No it was :\[1-\frac{ 1 }{ 4 }(2-\frac{ 3-x }{ 5 })\le \frac{ 8-x }{ 3 }\]

hero (hero):

I tried to get you to post something sooner.

OpenStudy (anonymous):

I am really sorry.

hero (hero):

\[1-\frac{ 1 }{ 4 }\left(2-\frac{ 3-x }{ 5 }\right)\le \frac{ 8-x }{ 3 }\] \[1-\frac{ 1 }{ 4 }\left(\frac{10}{5}-\frac{ 3-x }{ 5 }\right)\le \frac{ 8-x }{ 3 }\] \[1-\frac{ 1 }{ 4 }\left(\frac{10-(3-x)}{5}\right)\le \frac{ 8-x }{ 3 }\] \[1-\frac{ 1 }{ 4 }\left(\frac{10-3+x}{5}\right)\le \frac{ 8-x }{ 3 }\] \[1-\frac{ 1 }{ 4 }\left(\frac{7+x}{5}\right)\le \frac{ 8-x }{ 3 }\] \[1-\frac{7+x}{20}\le \frac{ 8-x }{ 3 }\] \[\frac{20}{20}-\frac{7+x}{20}\le \frac{ 8-x }{ 3 }\] \[\frac{20-(7+x)}{20}\le \frac{ 8-x }{ 3 }\] \[\frac{13-x}{20}\le \frac{ 8-x }{ 3 }\]

OpenStudy (anonymous):

Thank you.

hero (hero):

I'm sure you know you're not done

OpenStudy (anonymous):

But if i have to find x ?

hero (hero):

Yes, you have to continue to isolate until you have isolated x

hero (hero):

The good thing is, we don't have to worry about any denominator restrictions

OpenStudy (anonymous):

O.K. Thank you again.

OpenStudy (anonymous):

I have a question.

OpenStudy (anonymous):

@Hero ?

hero (hero):

What's the question

OpenStudy (anonymous):

If I substract the right side does it becomes : \[\frac{ 13-x }{ 20 }- \frac{ 8-x }{ 3 }\le0\] ?

hero (hero):

Yes, you could do that, but you don't have to

OpenStudy (anonymous):

I do not know if i am going to find x by this way.

OpenStudy (anonymous):

The x dissapears. I think.

hero (hero):

I take that back. There IS a solution. Continue solving for x.

hero (hero):

Just try solving for x

OpenStudy (anonymous):

I have to substract ?

hero (hero):

You could have simply cross multiplied

hero (hero):

\[\frac{13-x}{20}\le \frac{ 8-x }{ 3 }\] At this point you could have done this: \[3(13 - x) \le 20(8 - x)\]

hero (hero):

Are you solving for x or are you stuck?

hero (hero):

If you used my approach you would distribute to get \[39 - 3x = 160 - 20x\] Then place like terms on the same side: \[20x - 3x = 160 - 39\] \[(20 - 3)x = 121\] \[17x = 121\]

hero (hero):

Oops, the equals should be \(\le\)

OpenStudy (anonymous):

I WAS. Thanks to you. Thank you again @Hero.

OpenStudy (anonymous):

I had not seen that you had cross multiplied. Perhaps if i had seen i would not been have stucked.